OpenStudy (anonymous):

how many moles are present in 1.51 x 10^23 molecules of MgCL2

OpenStudy (anonymous):

Do you know the formula for calculating no. of moles in a given compound?

OpenStudy (anonymous):

well yes, but I think one of my steps is wrong

OpenStudy (australopithecus):

use avagadros number

OpenStudy (anonymous):

I found the molar mass of MgCl2 = 95.205 but I kind of guessed to mutiply the 1.51 x10^23 by 6.022 x10^23 then by 95.205 which was wrong.

OpenStudy (anonymous):

How do you find mole from mass and molar mass?

OpenStudy (anonymous):

Molar mass is correct.

OpenStudy (anonymous):

Yes the second guess is wrong.

OpenStudy (anonymous):

Let me think of this formula : \[\text{No. of moles} = \frac{\text{Mass in grams}}{\text{Molecular Mass}}\]

OpenStudy (australopithecus):

avagadros number = 6.02214129(27)×10^23/mol 1.51 x 10^23 molecules / 6.02214129(27)×10^23 molecules/mol

OpenStudy (anonymous):

Can you calculate this,,,,@Lanik what @Australopithecus had given ... She has made your way easier..

OpenStudy (australopithecus):

sorry avagadros number is 6.022141*10^23

OpenStudy (australopithecus):

but yeah it usually has unites of 1/mol

OpenStudy (anonymous):

Why you wrote ... (27) there ?

OpenStudy (australopithecus):

if you divide by it, it will convert molecules to mole

OpenStudy (australopithecus):

I cut and pasted it from wikipedia

OpenStudy (anonymous):

lol got it :)

OpenStudy (anonymous):

Cut? it must be copied ;)

OpenStudy (anonymous):

okay i got 2.50747 x 10^45

OpenStudy (anonymous):

Calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright ...

OpenStudy (anonymous):

So where are the units? Can you name the unit of that.?

OpenStudy (anonymous):

Thats a wrong answer. doesnt match. SOOO….

OpenStudy (anonymous):

Was it the final step...?

OpenStudy (australopithecus):

1.51 x 10^23/6.022141*10^23 = 0.25mol

OpenStudy (anonymous):

You have to use 95.05 g also.. So the above step was not the "last" step

OpenStudy (anonymous):

10^{23} will get cut..... from where did you get 10^45 lanik?

OpenStudy (anonymous):

thats so werid I did the exact same thing and I got the other answer

OpenStudy (anonymous):

 What is the answer given @Lanik (I mean the answer in the book)